OpenStudy (anonymous):

Evaluate the function rule to find the range of each function for the domain {-3, 0, 5}. y = -2x + 6 A. {-6, -4, 0} B. {-4, 0, 6} C. {-4, 6, 12} D. {6, 12, 16}

OpenStudy (jango_in_dtown):

y=f(x)=-2x+6 x take values -3,0,5 f(-3)=-2(-3)+6=6+6=12 f(0)=-2(0)+6=0+6=6 f(5)=-2(5)+6=-10+6=-4 so range={12,6,-4}={-4,6,12}